Suppose a golfer hits a ball with a velocity of 40 m/s at an angle of 30 degrees to the horizontal. The projectile model can be used to answer some questions about what will happen to the ball later during its flight.

a) where will the ball be 2 seconds later?

b) what is the greatest height the ball will reach?

c) if the golf ball is hit in the direction of a 12 meter tree which is 80 meters from the golfer, will the ball pass over the tree or hit it?
Answer #1

usin 30 : 1130 tol | vººdo a) veretical velocity: visin30*8 ms 2 400 m/= 20m/s. horizontal velocity 2 V.49s 300 ml 40x V3 - 3its height =CH-0.3531) m 20:0546m which shows that it will pass over the tree